Find the linearization of the function f(x) = 5 + x2 at x = 2. Explain how you arrive at your answer.
Answer #1

Consider the given function f(x) = √5 +2² then f(2)=√5+(2) = √5+4 ra and f(x) = then f) = 2 √5+02 √5+(2)² 3 2 33 The lineari